### v1

# The power-spectrum tensor in steady-state systems and its role in quantum friction

Abstract
For systems in equilibrium, quantum statistical physics provides a number of general theorems and relations that are not tied to specific microscopic models, one example being the fluctuation-dissipation theorem. Much less is known for non-equilibrium situations. In this work, we discuss certain properties of the power-spectrum tensor for systems in general steady-states, i.e. stationary states not necessarily corresponding to equilibrium configurations. In our analyses, we do not make any direct connection to specific models for the underlying microscopic dynamics and, therefore, our results can be applied to a large variety of systems. We also connect the power-spectrum tensor to other quantities that characterize these systems and, where appropriate, compare with the equilibrium counterparts. As an application, we consider the specific problem of quantum friction, where, at zero temperature, a contactless quantum-electrodynamic drag force acts on a particle that moves in close proximity to an arrangement of material bodies. Specifically, we show how the additional information about the system’s physics facilitates the derivation of more precise constraints on the power spectrum and its functional dependencies.
I Introduction
Stochastic processes are omnipresent in engineering, physics and chemistry, so that a detailed understanding of their properties lies at the heart of many applications both, for fundamental and technological purposes. Specifically, the power spectral density is one of the most relevant quantities for studying noisy time-stationary systems: It describes how the strength (commonly the power) of different components of stochastic processes is distributed across different frequencies [1]. In stationary systems it is connected to the correlation function via the Wiener-Khinchin theorem [1], establishing that one quantity is the Fourier transform of the other. Whenever the stochastic process is multivariate or multidimensional, both the correlation and the power spectral density become tensorial in nature and we speak of correlation and power-spectrum tensors, respectively. In addition to the autocorrelations of each of the elements of the stochastic process, these quantities also contain the cross-correlations between them.
In thermal equilibrium, a system’s power-spectrum tensor is connected to the system’s susceptibility tensor by way of the fluctuation-dissipation theorem (FDT) [2, 3]. The FDT is related to the detailed balance between the different energy-exchange processes that take place in equilibrium systems [4, 5]. Probably one of the most remarkable feature of this relation is its generality: As long as the system is in thermal equilibrium, the details of the underlying microscopic dynamics do not matter – whether the dynamics is linear or nonlinear, whether it is quantum or classical, whether it is Markovian or not [4, 1, 6, 7], the power spectrum is always proportional to the susceptibility via a known factor that depends on temperature. Therefore, it is not at all surprising that the FDT is applied in numerous and very diverse physical systems [8, 9, 10, 11, 12, 13, 14], sometimes even transcending its range of applicability. In fact, the FDT has been considered also for nonequilibrium systems for which, technically speaking, it becomes inaccurate [15, 16, 17, 18, 19, 20, 21, 22, 23]. Arguments for the validity of this approach have typically been based on the so-called local thermal equilibrium (LTE) approximation. Within the LTE approximation, it is assumed that the correlation length associated with the fluctuating dynamics is so short that the system’s spatially separated components are (locally) in thermal equilibrium with their immediate surroundings [24, 25, 26]. However, the quality of this approximation depends strongly on the specific system and its parameters, thus rendering it at least questionable in a number of cases [24, 26].
The desire for deeper insights into nonequilibrium physics, also in connection to the hindsight provided by stochastic thermodynamics [27, 28], has recently motivated the research on more general fluctuation-dissipation and thermodynamic uncertainty relations [6, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], i.e. connections between different physical quantities and bounding inequalities with a broad range of validity, even beyond equilibrium. Within this context, nonequilbrium and the breaking of equilibrium relations has been investigated theoretically [39, 40, 41, 42, 43] and experimentally in systems such as polymers [44], colloidal gases [45], and spin glasses [46]. Particular attention has been devoted to systems in nonequilibrium steady-states (NESS) [47, 41, 5, 48, 49, 50, 51, 52, 30]. In such stationary states, a system’s stochastic properties are translation invariant with respect to time, despite the fact that the system is not in equilibrium. To exist, the NESS must thus be sustained by one or several external agents such as externally maintained temperature gradients, chemical potential differences, or even mechanical drives that act on the system. In general, this can break the aforementioned detailed balance and, therefore, lead to violations of the FDT [5].
In this work, we also focus on a stationarity and discuss some properties of the power-spectrum tensor for a system that is characterized by being in a steady-state. While we address nonequilibrium dynamics, our approach also includes the equilibrium case as a special limit. Specifically, we do not make any detailed assumption about the microscopic dynamics of the system, thereby keeping the level of generality similar to that of the FDT. The stationarity of the NESS still allows to make certain general statements that can be related to the properties of the underlying stochastic dynamics.
We further discuss the properties of the power-spectrum tensor in the case of the nonconservative viscous motion of a particle in the quantum electromagnetic vacuum [53, 54]. Specifically, we address quantum friction, i.e. the drag force acting on a neutral particle driven with constant velocity in the empty space close to a generic arrangement of bodies. Still keeping the microscopic dynamics unspecified, we show how additional information about the physics of the system strongly constrains the functional behavior and the structure of the power-spectrum tensor.
The paper is organized as follows. In Sec. II, we consider the power-spectral density in the steady-state of a multidimensional stochastic process. In Sec. III, we specify our analysis to the case of quantum friction and show how the thus increased information about the system allows us to derive additional relations. To confirm our results and illustrate the advantages as well as the simplifications introduced by our general approach, we consider here an exactly solvable model for describing the internal dynamics of the particle. Finally, in Sec.IV we summarize and discuss our findings.
II The power-spectrum tensor in steady-state systems
Let us consider a generic physical system described by the Hamiltonian $\hat{H}_{0}$. Our focus here is on a quantum system with a large number of degrees of freedom. It is thought, in general, as being composed of a collection of interacting subsystems: On its own each of the quantum subsystems is open, i.e. able to exchange energy and other physical quantities with the remaining parts of the system, representing its environment [55, 56]. Contrary to typical approaches, however, we will refrain from deriving any master equation, avoiding the Markov and other approximations which, depending on the specific circumstances, can be problematic for the description of quantum systems [4, 7, 57]. Indeed, except for the fact that the full system is in a unique stationary state described by the density matrix $\hat{\rho}$ with $[\hat{\rho},\hat{H}_{0}]=0$, we do not rely on any specific microscopic information about its dynamics. Again, we would like to note explicitly that $\hat{\rho}$ does not necessarily correspond to an equilibrium state.
Let $\hat{d}_{i}$ be the components of a Hermitian vector-operator $\hat{\mathbf{d}}$. The time-evolution of each component, $\hat{d}_{i}(t)$, is induced according to $\hat{H}_{0}$. For our purposes, we consider a three-dimensional vector but generalizations to different dimensions are possible. The vector-operator $\hat{\mathbf{d}}$ either corresponds to a three-dimensional observable ($i=x,y,z$, for instance a particle’s electric dipole moment – see Sec. III), or a quantity that characterizes a multivariate stochastic process as for example in three-particle systems. For simplicity, in the subsequent discussion, we will refer to the former case but most of the results can be transferred to the latter case upon adequate replacements of the relevant quantities.
Since the system is stationary, the expectation value of $\hat{\mathbf{d}}(t)$ over $\hat{\rho}$ is constant. Without loss of generality, we assume it to be zero [58]. i.e. $\langle\hat{d}_{i}(t)\rangle\equiv\mathrm{tr}[\hat{d}_{i}(t)\hat{\rho}]=\mathrm{tr}[\hat{d}_{i}\hat{\rho}]=0$, where ‘$\mathrm{tr}$’ indicates the trace over a generic set of basis vector-states of the system’s Hilbert space. Given $\hat{d}_{i}(t)$ and $\hat{d}_{j}(t^{\prime})$, one can define the corresponding correlation function
$C_{ij}(\tau)\equiv\langle\hat{d}_{i}(t)\hat{d}_{j}(t^{\prime})\rangle=\mathrm{tr}\left[\hat{d}_{i}(t)\hat{d}_{j}(t^{\prime})\hat{\rho}\right]~{},$ (1)
which for a stationary state depends only on the time difference $\tau=t-t^{\prime}$. Owing to the fact that, in general, the involved operators do not commute, we may have complex-valued correlation functions even if the operators $\hat{d}_{i}$ are Hermitian [59] From Eq. (1) and the stationarity, we then further obtain that $C_{ij}^{*}(\tau)=C_{ji}(-\tau)$, where the subscript ‘*’ denotes complex conjugation. The set of correlation functions represents a correlation tensor $\underline{C}(\tau)$ so that we may introduce a matrix notation for which we accordingly have that
$\underline{C}^{{\dagger}}(\tau)=\underline{C}(-\tau).$ (2)
Here, the superscript $\dagger$ denotes the Hermitian conjugate of a matrix.
Using the Wiener-Khinchin theorem, we define the power-spectrum tensor $\underline{S}(\omega)$ as
$\underline{S}(\omega)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\tau}{2\pi}\;\underline{C}(\tau)\;e^{\mathrm{i}\omega\tau}.$ (3)
We start the analysis of this quantity by noticing that Eq. (2) implies $\underline{S}^{{\dagger}}=\underline{S}$ so that the power spectrum is necessarily Hermitian. As a consequence, we can always write
$\underline{S}(\omega)=\underline{K}(\omega)+\bm{\Omega}(\omega)\cdot\mathbf{\underline{L}}~{},$ (4)
where $\underline{K}(\omega)$ is a real symmetric tensor. The quantity $\bm{\Omega}(\omega)$ is a real three-dimensional vector so that $\bm{\Omega}\cdot\mathbf{\underline{L}}\equiv\sum_{k}\Omega_{k}\cdot\underline{L}_{k}$ with $[\underline{L}_{k}]_{ij}=-\mathrm{i}\epsilon_{kij}$ being the skew-symmetric (trace-less) matrix generators of the real group of three-dimensional rotations (SO(3) group; $\epsilon_{kij}$ indicates the Levi-Civita symbol).
The first and second term on the r.h.s. of Eq. (4) correspond, respectively, to the real and imaginary part of the power-spectrum tensor, thereby uniquely identifying their associated symmetry properties. One can show the power-spectrum tensor $\underline{S}(\omega)$ is a positive semidefinite matrix (see App. A). Consequently $\underline{K}(\omega)=[\underline{S}(\omega)+\underline{S}^{\sf T}(\omega)]/2$ (the superscript ‘$\sf T$’ denotes the transpose of a matrix) is also positive semidefinite, since it is given by the sum of two positive semidefinite matrices. Given that the orthogonal transformation that diagonalizes $\underline{K}(\omega)$ is an element of SO(3), we can find a basis where in the decomposition of Eq. (4) $\underline{K}(\omega)$ is diagonal. We can write the vector $\bm{\Omega}(\omega)$ as the Fourier transform of
$\bm{\Omega}(\tau)=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(t)\times\hat{\mathbf{d}}(t^{\prime})\rangle=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(\tau)\times\hat{\mathbf{d}}(0)\rangle$ (5)
for which we obtain $\bm{\Omega}(-\tau)=\bm{\Omega}^{*}(\tau)$. One can interpret this as a measure for the net rotation of $\hat{\mathbf{d}}(t)$ with respect to $\hat{\mathbf{d}}(t^{\prime})$. For classical systems, a similar quantity has been associated with the occurrence of nonequilibrium dynamics [60, 61, 62], pointing to an analogous meaning for part of $\underline{S}(\omega)$ proportional to $\bm{\Omega}(\omega)$.
In connection with the power-spectrum tensor, it is also convenient to study the susceptibility tensor for the same stationary system. It is defined as the Fourier transform of the corresponding linear response tensor $\underline{\alpha}(\tau)$. For our three-dimensional vectorial quantum-mechanical observable, the elements of the response tensor are given by (see App. A)
$\alpha_{ij}(t-t^{\prime})\equiv\theta(t-t^{\prime})\langle\frac{\mathrm{i}}{\hbar}[\hat{d}_{i}(t),\hat{d}_{j}(t^{\prime})]\rangle~{},$ (6)
where $\theta(t)$ denotes the Heaviside step function Once again, due to stationarity, the linear response depends only on $\tau=t-t^{\prime}$. Given that $\underline{\alpha}(\tau)$ is real, its Fourier transform fulfills what we call ‘crossing relation’, i.e. $\underline{\alpha}^{*}(\omega)=\underline{\alpha}(-\omega)$.
As long as the Fourier transform exists, the susceptibility $\underline{\alpha}(\omega)$ is analytical in the upper half of the complex frequency plane. As a consequence, it fulfills the following properties: Along the positive imaginary frequency axis, $\alpha_{ij}(\omega)$ is a real and monotonically decreasing function [63] and its real and imaginary part are, respectively, even and odd functions of $\omega$. The exchange (absorption and emission) of energy connected with our vectorial quantum-mechanical observable during a generic perturbation can be written in terms of
$\underline{\alpha}_{\Im}(\omega)=\frac{\underline{\alpha}(\omega)-\underline{\alpha}^{{\dagger}}(\omega)}{2\mathrm{i}}~{},$ (7)
which is proportional to the skew-Hermitian part of $\underline{\alpha}(\omega)$. Contrary to the linear susceptibility tensor, for $\omega>0$ $\underline{\alpha}_{\Im}$ is a positive definite Hermitian matrix (see App. A) and, similar to $\underline{S}$, allows for a decomposition analogous to Eq. (4). Owing to the crossing relation, we have that $\underline{\alpha}_{\Im}(-\omega)=-\underline{\alpha}^{*}_{\Im}(\omega)=-\underline{\alpha}^{\sf T}_{\Im}(\omega)$, which implies that the real part of each matrix element is odd in $\omega$, while the corresponding imaginary part is even in the same variable. As a consequence, we also have that in general $\underline{\alpha}_{\Im}(0)\neq 0$ and is skew-symmetric. The elements of $\underline{\alpha}_{\Im}(\omega)$ are directly related to the commutator appearing in Eq. (6). In particular, we obtain
$\langle[\hat{d}_{i}(t),\hat{d}_{j}(t^{\prime})]\rangle=\frac{\hbar}{\pi}\int_{-\infty}^{\infty}\mathrm{d}\omega\;[\underline{\alpha}_{\Im}(\omega)]_{ij}\;e^{-\mathrm{i}\omega(t-t^{\prime})}~{},$ (8)
which can also be interpreted as a consequence of the analytical properties of the susceptibility tensor.
For the following considerations, it is convenient to introduce the matrix $\underline{\nu}(\omega)$ according to
$\underline{\nu}(\omega)=\underline{S}(\omega)-\frac{\hbar}{2\pi}\underline{\alpha}_{\Im}(\omega)~{}.$ (9)
This matrix is Hermitian [64, 65, 32], positive semidefinite (see App. A) and obeys the crossing relation $\underline{\nu}(\omega)=\underline{\nu}^{*}(-\omega)$. Its elements are the Fourier transforms of $\nu_{ij}(\tau)=\langle\{\hat{d}_{i}(t),\hat{d}_{j}(t^{\prime})\}\rangle/2$, which are real-valued as they represent the expectation values of the anti-commutators of Hermitian operators. Consequently, the real and imaginary parts of $\nu_{ij}(\omega)$ are, respectively, even and odd functions of $\omega$. In fact, the matrix $\underline{\nu}(\tau)$ is the symmetrized version [64, 65] of the correlation tensor defined in Eq. (1). From Eq. (9), taking the difference between $\underline{\nu}(\omega)$ and $\underline{\nu}^{*}(-\omega)$ and using the properties of the involved matrices, we find
$\underline{S}^{*}(-\omega)=\underline{S}(\omega)-\frac{\hbar}{\pi}\underline{\alpha}_{\Im}(\omega)~{}.$ (10)
Since $\underline{S}$ and $\underline{\alpha}_{\Im}$ are, respectively, positive semidefinite and positive for $\omega>0$, we obtain that $\underline{S}(\omega)>\underline{S}^{*}(-\omega)=\underline{S}^{\sf T}(-\omega)$, where the inequality has to be interpreted in the sense of the Loewner order [66]. The asymmetry of Eq. (10) is shaped by the quantum nature of the system. Indeed, classically one has that $\underline{C}(\tau)\in\mathbb{R}$ implying that $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$.
In the special case of a system in thermal equilibrium at the temperature $T$, i.e for the specific steady state given by
$\hat{\rho}=\frac{e^{-\beta\hat{H}_{0}}}{Z},\quad Z=\mathrm{tr}[e^{-\beta\hat{H}_{0}}]$ (11)
where $\beta=1/[k_{B}T]$, the power spectrum and the susceptibility are connected by the expression
[…]
Indeed, this one of of possible forms of the celebrated fluctuation-dissipation theorem (FDT) [2, 3]. Given that equilibrium is a special steady state, we can verify that Eq. (12) must explicitly fulfill Eq. (10). Specifically, here Eq.(10) can be seen as the manifestation of the asymmetry introduced by quantum fluctuations in the absorption and emission processes characterizing the system dynamics. A concrete physical example of this phenomenon is the spontaneous decay process characterizing an atom in the electromagnetic vacuum [67].
In equilibrium, the Onsager reciprocity relations [68, 69, 63] imply that the susceptibility $\underline{\alpha}$ must be symmetric so that $\underline{\alpha}_{\Im}=\underline{\alpha}_{I}$, where the subscript ‘$I$’ indicates the element-wise imaginary party of the tensor. As a consequence, the corresponding power-spectrum tensor is real and symmetric. The Onsager reciprocity relations are thus equivalent to the condition $\bm{\Omega}=0$ which indicates that in equilibrium, despite possible (quantum) fluctuations, no net rotation of $\hat{\mathbf{d}}(\tau)$ around $\hat{\mathbf{d}}(0)$ can occur [60, 61, 62]. Equivalently, from Eq. (2), we have that the correlation tensor must fulfill the crossing relation (in time) $\underline{C}^{*}(\tau)=\underline{C}(-\tau)$ (the correlation tensor becomes symmetric).
For an arbitrary steady-state, we now write a relation which resembles Eq. (12):
[…]
Here, $\beta$ corresponds to the local equilibrium temperature of the system, often coinciding with that of its immediate surrounding. The matrix $\underline{J}(\omega)$ represents the correction to the FDT due to the system’s possible nonequilibrium dynamics. In fact, we can relate this correction to the net energy that flows between the system and the environment in the steady-state [70, 71, 72, 73, 36, 26].
The definition in Eq. (13) allows us to render the question about the validity of the local thermal equilibrium approximation more precise, i.e. as a quantitative comparison of the impact of $\underline{J}(\omega)$ on the relevant physical quantities relative to that of the FDT-like term. As long as the contribution of $\underline{J}(\omega)$ can be neglected, the local thermal equilibrium approximation provides a viable approach. For a more detailed analysis, one often has to rely on the system’s microscopic dynamics and on the specific physical quantity of interest (see also Sec. III.2). Nonetheless, we can infer a number of general properties of $\underline{J}(\omega)$. Indeed, since both $\underline{S}(\omega)$ and $\underline{\alpha}_{\Im}(\omega)$ are Hermitian, we must have that $\underline{J}(\omega)=\underline{J}^{{\dagger}}(\omega)$. Upon employing the definition of $\underline{\nu}(\omega)$, we can write
$\underline{J}(\omega)=\underline{\nu}(\omega)-\frac{\hbar}{2\pi}\coth\left(\frac{\beta\hbar\omega}{2}\right)\underline{\alpha}_{\Im}(\omega)~{}.$ (14)
Therefore, we can conclude that $\underline{J}(\omega)$ fulfills the crossing relation $\underline{J}(\omega)=\underline{J}^{*}(-\omega)=\underline{J}^{\sf T}(-\omega)$, highlighting that $\underline{J}(\tau)$ is real. The vanishing of $\underline{J}(\omega)$ provides yet another version of the (equilibrium) FDT while $\underline{J}(\omega)\neq 0$ gives a measure of the violation of the detailed balance within NESS [70, 71, 72, 73, 26]. Specifically, for $T=0$ and $\omega>0$ we obtain
$\underline{J}(\omega)=\underline{\nu}(\omega)-\frac{\hbar}{2\pi}\underline{\alpha}_{\Im}(\omega)=\underline{S}^{*}(-\omega)~{},$ (15)
which allows us to state that in the limit of vanishing temperature $\underline{J}(\omega)$ must be positive semidefinite. In the same limit, we can write
$J_{ij}(\tau)=$ $\frac{1}{4\pi}\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle$
$-\frac{1}{(2\pi)^{2}}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}\tau^{\prime}\;\frac{\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle}{\tau^{\prime}-\tau}~{},$ (16)
which specifically demonstrates how the conditions for the validity of the FDT imply the vanishing of $\underline{J}$. Indeed, in equilibrium, the correlations $C_{ij}(\tau)=\langle\hat{d}_{i}(\tau)\hat{d}_{j}(0)\rangle=\langle\{\hat{d}_{i}(\tau),\hat{d}_{j}(0)\}\rangle/2-\mathrm{i}\langle\mathrm{i}[\hat{d}_{i}(\tau^{\prime}),\hat{d}_{j}(0)]\rangle/2$ become analytical so that their real and imaginary part must fulfill the Kramer-Kronig relations [74], for which the terms on the r.h.s. of Eq. (II) cancel to zero. More precisely, we would like to recall that the correlation tensor $\underline{C}(\tau)$ in the equilibrium state at a generic temperature $T$ satisfies the following two additional properties: Firstly, the elements of the matrix $\underline{C}^{*}(\tau)$ are analytical functions within the stripe $0<\tau_{I}<\hbar\beta$ in the upper half part of the complex $\tau$-plane [75]; Secondly, in an equilibrium state $\underline{C}(\tau)$ fulfills the so-called Kubo-Martin-Schwinger condition, i.e. $\underline{C}^{*}(\tau)=\underline{C}(\tau-\mathrm{i}\hbar\beta)$  [76, 3, 75, 74, 55].
The previous considerations illustrate that, in addition to possibly having $\bm{\Omega}\neq 0$, the system’s quantum nonequilibrium dynamics is in general characterized by $\underline{J}\neq 0$. Although both quantities are connected to the breaking of the detailed balance occurring at equilibrium, the underlying concepts they are addressing are, in general, not identical. As we will see below, there are indeed situations for which $\bm{\Omega}=0$ and $\underline{J}\neq 0$.
III Viscous motion through the electromagnetic vacuum
As soon as the system becomes more concrete, its specific properties and symmetries allow us to draw further conclusions on its stochastic dynamics. In this section we illustrate this aspect within the framework of quantum electrodynamics, where the quantum fluctuations pervading the system are those connected with the absorptive, dissipative and dispersive processes characterizing light-matter interaction. As a specific example, we consider a electrically neutral particle (e.g. an atom) which moves in vacuum with nonrelativistic velocity along a prescribed trajectory. We further assume that the trajectory of its center of mass is classical (usually a good approximation for particles with relatively large masses or velocities) and parallel to the translationally invariant direction of an arbitrary arrangement of electrically neutral material bodies, see Fig. 1. We denote the direction of motion with the unit vector $\mathbf{n}$ and we use the vector $\mathbf{R}_{a}$ to identify the particle’s lateral position. The material bodies are comprised of linear, reciprocal and passive materials which we further assume to be nonmagnetic, dissipative, possibly spatially dispersive and in thermal equilibrium with the surrounding electromagnetic field.
Figure 1: Schematic representation of a particle moving in vacuum close to an arrangement of different, macroscopic objects. The particle moves on a straight line with direction $\mathbf{n}$, which coincides with the direction of translational invariance of the arrangement of macroscopic bodies. In this setup, the quantum fluctuations of the electromagnetic field behave as a viscous medium and the particle’s motion is affected by a corresponding drag force. An external force acting on the particle’s center of mass (not shown) keeps the motion at constant velocity $v$.
Despite the system being electrically neutral, when the particle moves along its trajectory, it interacts with the quantum fluctuations of the electromagnetic field which, by the presence of the material bodies, are modified relative to vacuum. As a result of this interaction a frictional force acts on the particle [77, 18, 53] (see also below) that tends bring the particle to rest with respect to the privileged reference frame defined by the arrangement of the material bodies. Consequently, an external force acting on the particle’s center of mass is required to maintain the particle’s motion. When an external drive and the frictional force balance, the system reaches a nonequilibrium steady-state characterized by the particle moving at constant speed $v$.
The electromagnetic viscosity experienced by the particle [54] is a manifestation of the fluctuating and dissipative nature of the system. Due to the existence of quantum fluctuations, it has the peculiarity to exist even at zero temperature, where it is usually referred to as quantum friction. Within the standard framework, the stationary fluctuating dynamics of the particle is usually described by its three-dimensional electric dipole (vector-)operator $\hat{\mathbf{d}}(t)$. In the above-discussed nonequilibrium steady-state, the corresponding correlation tensor $\underline{C}(\tau,v)$ and the power-spectrum tensor $\underline{S}(\omega,v)$ explicitly depend on the particle velocity $v$ which thus plays the role of a control variable of our nonequilibrium configuration [33]. Naturally, $v=0$ indicates the equilibrium state. In general, these tensors also depend on the position $\mathbf{R}_{a}$ but in order to avoid notational clutter, in the subsequent discussions, we suppress the dependence on $\mathbf{R}_{a}$ unless explicitly required. Clearly, for any value of $v$ and $\mathbf{R}_{a}$ the power-spectrum tensor and the related quantities must fulfill the relations discussed in Sec. II.
In the NESS, the particle’s dipole does not only viscously move with velocity $v$ but, due to the interaction with its quantum electromagnetic environment, it can also gain and loose angular momentum [25, 26, 78]. However, owing to the system’s symmetry and its intrinsic dissipative behavior, in the stationary state the particle cannot have a net angular momentum with the direction parallel to $\mathbf{n}$. As a physical consequence, the dipole cannot feature a net averaged rotation around the direction of motion and Eq. (4) takes the form
$\underline{S}(\omega,v)=\underline{K}(\omega,v)+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~{},$ (17)
where $\perp$ indicates a direction orthogonal to $\mathbf{n}$.
Equation (17) can be used as starting point for discussing some general aspects about the tensorial structure of $\underline{S}(\omega,v)$ and its functional dependence on $v$. Indeed, given that for $v=0$ the system is in equilibrium and, therefore, the power spectrum is real and symmetric, we can already conclude that $\bm{\Omega}_{\perp}(\omega,0)=0$. Since the consituent materials are reciprocal, we have
$\underline{S}(\omega,v)=\underline{M}_{n}\underline{S}(\omega,-v)\underline{M}^{-1}_{n}~{}.$ (18)
Here, we denote with $\underline{M}_{n}=\underline{\mathds{1}}-2\mathbf{n}\mathbf{n}^{\sf T}$ the transformation that describes the mirror symmetry with respect to the plane orthogonal to $\mathbf{n}$ ($\underline{\mathds{1}}$ is the identity matrix). The change in sign of the velocity results from applying $\underline{M}_{n}$ to the velocity vector $\mathbf{v}=v\mathbf{n}$, i.e. $\underline{M}_{n}\mathbf{v}=-v\mathbf{n}$. However, the same transformation leaves $\mathbf{R}_{a}$ and all (proper) vectors parallel to the plane of symmetry unchanged. Further, we find it convenient to decompose the real matrix $\underline{K}$ as
$\underline{K}(\omega,v)=\underline{\Delta}(\omega,v)+\bm{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}~{},$ (19)
where $\underline{\Delta}$ is diagonal and $\bm{\chi}$ is a three-dimensional real-valued vector. The matrices $[\underline{\Lambda}_{k}]_{ij}=|\epsilon_{kij}|$ are symmetric, traceless and trace-orthogonal to $\underline{L}_{i}$ and $\underline{\Delta}$. In particular, due to their connection with the Gell-Mann matrices [79], the decomposition over all the previous matrices is unique.
If we denote with $\underline{\Lambda}_{n}=\mathbf{n}\cdot\mathbf{\underline{\Lambda}}$ the matrix that corresponds to the direction $\mathbf{n}$, we obtain that $\underline{\Lambda}_{n}$ and $\underline{\Delta}$ are invariant with respect to $\underline{M}_{n}$. For the remaining matrices, we find the transformations $\underline{M}_{n}\bm{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\bm{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. Due to the uniqueness of the decomposition, we can thus conclude from Eq. (18) that $\underline{\Delta}(\omega,v)$ and $\chi_{n}(\omega,v)$ are even in the velocity $v$, while $\bm{\chi}_{\perp}(\omega,v)$ and $\bm{\Omega}_{\perp}(\omega,v)$ are odd functions of $v$. Taking into account that $\underline{L}^{\sf T}_{i}=-\underline{L}_{i}$ we can further write (see App. B)
$\underline{S}(\omega,-v)=\underline{S}^{\sf T}(\omega,v)-2\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~{},$ (20)
which represents the formal counterpart of Eq. (10) for the variable $v$. By proceeding in an analogous way or by exploiting their connection to the power-spectrum tensor, we can derive similar relations for the matrices discussed in the previous section. Specifically, from Eq. (10) we obtain
$\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)-2\pi\frac{\bm{\chi}_{\perp}(\omega,v)-\bm{\chi}_{\perp}(-\omega,v)}{\hbar}\cdot\mathbf{\underline{\Lambda}}_{\perp}~{}.$ (21)
An interesting special case for this relation occurs if the expression for the susceptibility of the quantum dipole operator is identical to that of its classical counterpart or does not depend on temperature. Under these circumstances, the last term of Eq. (21) must always coincide with its classical limit. Since we already know that, classically, $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$, we directly infer that $\bm{\chi}_{\perp}(-\omega,v)=\bm{\chi}_{\perp}(\omega,v)$. Therefore, the last term of Eq. (21) vanishes, leading to the identity $\underline{\alpha}_{\Im}(\omega,-v)=\underline{\alpha}_{\Im}^{\sf T}(\omega,v)$. Due to its analyticity and the Kramers-Kronig relations, this means that when the susceptibility has the same form within the quantum and classical description, it must fulfill $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$. This is tantamount to stating that, in this case, the system fulfills Onsager–Casimir-like relations where the atom’s speed $v$ takes the role of the external bias parameter [80].
If the system exhibits an additional mirror-symmetry $\underline{M}_{t}$, where $\mathbf{t}$ is a unit vector orthogonal to $\mathbf{n}$, then we can show along the lines discussed above that $\bm{\chi}=\chi_{t}\mathbf{t}$ and $\bm{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$ (see App. B). However, while the symmetry described by $\underline{M}_{n}$ is independent of the transverse location $\mathbf{R}_{a}$, the symmetry associated with $\underline{M}_{t}$ can, in general, only exist when $\mathbf{R}_{a}$ is located within (or parallel to) the symmetry plane. In other words, conclusions about $\underline{M}_{t}$ do depend on the transverse position of the particle. Similarly, if an additional mirror-symmetry along the direction $\mathbf{n}\times\mathbf{t}$ exists, this necessarily implies that, when the particle moves along the associated symmetry axis (located at the intersection of the two corresponding symmetry planes), the power-spectral tensor is diagonal, suppressing any possible rotation of the dipole operator  [25, 78, 54]. As mentioned above, although in this case the particle is characterized by a specific position $\mathbf{R}_{a}$, this is an example of nonequilibrium dynamics for which $\bm{\Omega}=0$. In addition, we would like to point out that the above reasoning essentially identifies a specific physical reference frame with one of the axes coinciding with $\mathbf{n}$, within which the power-spectrum tensor has a diagonal form. Indeed, while a unitary transformation that diagonalizes the power-spectrum tensor always exists, this transformation does, in general, not correspond to a physical rotation.
So far, we have assumed that the three components of the dipole operator can be varying in time independent of each other. Some of the above conclusions are modified for a particle’s model in which the dynamics along the three axes is correlated by construction. This occurs, for instance, if we consider a common model in quantum optics where the particle’s dipole operator is modeled as $\hat{\mathbf{d}}=\mathbf{d}\hat{\sigma}$, where $\mathbf{d}$ is a constant three-dimensional real-valued vector and $\hat{\sigma}$ is a scalar operator – usually a Pauli operator for a simple two-state system [81, 24, 82, 83, 84] or a 1D-harmonic oscillator [57, 85]. In this case, the correlation tensor is always proportional to the real and symmetric dyadic $\mathbf{d}\mathbf{d}^{\sf T}$, i.e.
$\underline{C}(\tau,v)=\mathbf{d}\mathbf{d}^{\sf T}\langle\hat{\sigma}(t+\tau)\hat{\sigma}(t+\tau)\rangle=\mathbf{d}\mathbf{d}^{\sf T}C_{\sigma}(\tau,v)~{}.$ (22)
As a consequence, we find $\underline{C}(\tau,v)=\underline{C}^{\sf T}(\tau,v)$ and from the general properties of the correlation function we obtain $C_{\sigma}(\tau,v)=C^{*}_{\sigma}(-\tau,v)$. The corresponding power-spectrum tensor $\underline{S}(\omega,v)=\mathbf{d}\mathbf{d}^{\sf T}\Sigma(\omega,v)$ is real and symmetric, implying that within this model $\bm{\Omega}$ vanishes exactly, irrespective of the particle’s position. From the physical point of view, this is equivalent to the inability of the dipole vector to rotate around its initial value, which is consistent with having the real vector $\mathbf{d}$ being time-independent. This means that, within this model, the particle cannot absorb or emit angular momentum  [25]. Using this model, we obtain from symmetry considerations that $\Sigma(\omega,v)$ must be an even function of $v$. We can see this most easily by considering the case where the dipole operator is oriented orthogonal to the direction of motion, i.e. $\mathbf{d}\equiv\mathbf{d}_{\perp}$. In this case, Eq. (18) in conjunction with fact that the matrix $\mathbf{d}_{\perp}\mathbf{d}_{\perp}^{\sf T}$ is invariant with respect to $\underline{M}_{n}$, we find that $\Sigma(\omega,v)=\Sigma(\omega,-v)$.
III.1 The power spectrum in quantum friction
In order to see how the previous information is relevant for understanding the physics of the system, let us consider the expression for the quantum frictional force. If the system is at zero temperature, earlier work  [73, 57, 25, 78] has shown that in the NESS the frictional force is oriented along the direction of translational invariance, i.e. $\mathbf{F}_{\rm fr}=F_{\rm fr}\mathbf{n}$ (see Fig. 1), and is of the form
$F_{\rm fr}=-2\int_{0}^{\infty}\mathrm{d}\omega\int\frac{\mathrm{d}q}{2\pi}\,q\;\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{-}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right].$ (23)
where we indicate with ‘$\mathrm{Tr}$’ the trace over the tensors’ indices, while $q$ is the component of the wave vector along the direction of motion and $\omega^{\pm}_{q}=\omega\pm qv$ is the Doppler-shifted frequency. In addition to the dipole’s power-spectrum tensor $\underline{S}$, the interaction is described in terms of the electromagnetic Green tensor $\underline{G}$: Physically, it represents the linear electromagnetic susceptibility tensor of the system without the particle. For the aforementioned arrangement of the material bodies that comprise the system, it can be in principle obtained by solving the corresponding Maxwell equations with a source term [57, 86].
The physical interpretation of Eq. (23) identifies the l.h.s. as the net total momentum per unit of time that is transferred to or from the particle during the stochastic dynamics connected with the underlying absorption and emission processes characterizing the interaction [87, 88, 25, 54]. Specifically, $\underline{G}_{\Im}=[\underline{G}-\underline{G}^{{\dagger}}]/(2\mathrm{i})$ in Eq. (23) is connected with the electromagnetic density of states [89]. Moreover, in complete analogy to $\alpha_{\Im}$, we may relate the matrix $\underline{G}_{\Im}$ to the exchange of energy (dissipation) with the electromagnetic environment that surrounds the particle and is consequently positive semidefinite for $\omega>0$ (see also the discussion around Eq. (30)).
Technically, the Green tensor represents a propagator of the electromagnetic field and as such usually depends on two spatial arguments [90, 91], a source position and a detection position. However, since in all our expressions we require the back-action of the electromagnetic field emitted by the particle onto itself (via scattering by the material bodies), source and detection positions coincide and we retain in the Green tensor only one spatial argument. In addition, given the system’s translational invariance along the particle trajectory, we carry out a Fourier transform, so that the Green tensor acquires a mixed Fourier- and real-space dependence with regards to the positional arguments and, of course, features a dependence on frequency, i.e. $\underline{G}(q,\mathbf{R}_{a},\omega)$. This expression fulfills the identities
$\underline{G}_{\Im}(-q,\mathbf{R}_{a},\omega)$ $=\underline{G}^{\sf T}_{\Im}(q,\mathbf{R}_{a},\omega)~{},$ (24a)
$\underline{G}_{\Im}(q,\mathbf{R}_{a},-\omega)$ $=-\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)~{},$ (24b)
which we may, respecitively, view as consequences of the system’s reciprocity and stationarity.
Furthermore, since $\underline{G}_{\Im}$ is Hermitian, we can decompose it into the sum of a diagonal matrix and a linear combination with real-valued coefficients of the matrices $\underline{L}_{i}$ and $\underline{\Lambda}_{i}$. From the reciprocity of the materials, given the mirror symmetry of the macroscopic bodies, we then have that
$\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)=\underline{M}_{n}\underline{G}_{\Im}(-q,\mathbf{R}_{a},\omega)\underline{M}^{-1}_{n}~{}.$ (25)
This would lead to an expression of the same form as provided in Eq. (20). However, due to the (system-specific) reciprocity condition, Eq. (24a), we can go one step further and find that the coefficients associated with $\bm{\underline{\Lambda}}_{\perp}$ and $\underline{L}_{n}$ actually vanish (see App. B) so that we can write
$\underline{G}_{\Im}=\underline{D}+c_{n}\underline{\Lambda}_{n}+\mathbf{s}_{\perp}\cdot\mathbf{\underline{L}}_{\perp}~{}.$ (26)
Here, the matrix $\underline{D}(q,\mathbf{R}_{a},\omega)$ is real-valued and diagonal. Both, the matrix $\underline{D}(q,\mathbf{R}_{a},\omega)$ and the function $c_{n}(q,\mathbf{R}_{a},\omega)$ are even in $q$. The real vector $\mathbf{s}_{\perp}(q,\mathbf{R}_{a},\omega)$ is odd in $q$ and can be related to a spin-dependent part of the electromagnetic density of states  [25, 92], thereby describing the so-called spin-momentum locking of light  [93, 94, 95, 96, 97]. Analogous to the power-spectrum tensor, if the arrangement of material bodies is also mirror-symmetric with respect to $\underline{M}_{t}$ and $\mathbf{R}_{a}$ lies in (or is parallel to) the plane of symmetry, we obtain $c_{n}=0$ and $\mathbf{s}_{\perp}=s_{t}\mathbf{t}$. By the same token, if a symmetry axis exist and $\mathbf{R}_{a}$ lies on this axis, $\underline{G}_{\Im}$ is diagonal.
Combining the results on $\underline{G}_{\Im}$ with those on the power-spectrum tensor, we can infer a number of properties of the frictional force. When the three components of the dipole operator can have independent dynamics and upon considering Eqs. (17), (19) and (26), we can write the trace in the expression for the frictional force as
$\mathrm{Tr}\left[\underline{S}^{\sf T}\underline{G}_{\Im}\right]=\mathrm{Tr}\left[\underline{\Delta}\underline{D}\right]+2\chi_{n}c_{n}-2\bm{\Omega}_{\perp}\cdot\mathbf{s}_{\perp}.$ (27a)
Here, we have utilized that $\mathrm{Tr}[\underline{L}_{i}\underline{L}_{j}]=\mathrm{Tr}[\underline{\Lambda}_{i}\underline{\Lambda}_{j}]=2\delta_{ij}$ and $\mathrm{Tr}[\underline{L}_{i}\underline{\Lambda}_{j}]=0$. For the case where $\hat{\mathbf{d}}(t)=\mathbf{d}\sigma(t)$ we find instead
$\mathrm{Tr}\left[\underline{S}^{\sf T}\underline{G}_{\Im}\right]$ $=\mathbf{d}\cdot\underline{G}_{\Im}\cdot\mathbf{d}\;\Sigma=\mathbf{d}\cdot\left(\underline{D}+c_{n}\Lambda_{n}\right)\cdot\mathbf{d}\;\Sigma$
=$\stackrel{{\scriptstyle\text{average}}}{{=}}\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{G}_{\Im}\right]\;\Sigma=\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{D}\right]\;\Sigma~{}.$ (27b)
The last line in the above equations corresponds to the result of an average over all possible directions of the vector $\mathbf{d}$ for which the power-spectrum tensor becomes proportional to $\underline{\mathds{1}}d^{2}/3$ ($d^{2}=\mathbf{d}\cdot\mathbf{d}$)  [24, 98, 82]. When comparing the two results, we notice, as expected, that the main difference lies in the contribution arising from $\bm{\Omega}_{\perp}$, which is absent in the second model. Moreover, when $\bm{\Omega}_{\perp}$ and $\mathbf{s}_{\perp}$ have the same sign, the corresponding contribution tends to reduce the frictional force. This confirms and generalizes earlier results obtained within a specific model for the dipole’s dynamics [25, 78] (see Sec. III.2). The remaining contributions in Eqs. (III.1) are similar in both models although not identical. Interestingly, the terms proportional to $c_{n}$ directly show how the system’s symmetry (or better the lack thereof) affects the frictional force. More precisely, as soon as the total system is mirror-symmetric with respect to a plane containing the particle’s trajectory – for example when the particle moves above a planar structure – the terms proportional to $c_{n}$ in Eqs. (III.1) vanish. Also, averaging over the dipole direction in the second model, washes out this contribution which does not appear in the last line of Eq.(III.1).
Another interesting consequence is that the results in Eqs. (III.1) are both equivalent to the identity
$\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{+}_{q},-v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]\\ =\mathrm{Tr}\left[\underline{S}(-\omega^{+}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]~{}.$ (28)
When used in Eq. (23) in conjunction with Eq. (24a) and after changing $q\to-q$, the previous expression allows us to show that, in general, the frictional force must always reverse its sign if $v\to-v$, i.e. $F_{\rm fr}(-v)=-F_{\rm fr}(v)$. In other words, independently from the model for the dipole’s dynamics, $F_{\rm fr}(v)$ is an odd function of the velocity. This is what is physically expected for a frictional interaction and has been confirmed by calculations with specific models [99, 100, 101, 73, 88, 54]. We want to underline, however, that the mathematical properties justifying this behavior are less trivial that one could intuitively presume. Indeed, reversing the logic and imposing that $F_{\rm fr}(v)$ is odd with respect to the velocity one would obtain Eq. (28), which can erroneously suggest that $\underline{S}^{\sf T}(\omega,-v)=\underline{S}(\omega,v)$, in contradiction with Eq. (20).
III.2 A illustrative example
So far, we have refrained from considering any specific model for the microscopic dynamics of the electric dipole (vector-)operator. In order to illustrate the validity and physical meaning of our results, we now proceeed to consider a specific model which allows for an exact solution of the equations of motions. Specifically, we assume that the particle moving in vacuum parallel to a translational invariant arrangement of bodies is an atom. Given that the interaction between the atomic internal degrees of freedom and the electromagnetic fluctuations is typically weak, it is reasonable to describe the dipole operator in term of linear system associated with a single frequency $\omega_{a}$. Physically, $\omega_{a}$ would correspond to the transition frequency between the highest occupied orbital to the lowest unoccupied one. Within this linear approximation, we are going to represent $\hat{\mathbf{d}}$ in terms of a 3D isotropic quantum-mechanical oscillator  [102, 73] which fulfills the following equation of motion
$\left(\partial_{t}^{2}+\omega_{a}^{2}\right)\hat{\mathbf{d}}(t)=\alpha_{0}\omega_{a}^{2}\hat{\mathbf{E}}(\mathbf{r}_{a}(t),t)~{}.$ (29)
In this expression, $\alpha_{0}$ is a constant which essentially assesses the strength of the coupling between the atom and the electromagnetic radiation. Specifically, it can be identified with the static polarizability associated with the aforementioned atomic transition  [57, 25].
The r.h.s. of Eq. (29) denotes the total fluctuating electric field that drives the dipole and it is evaluated at the position $\mathbf{r}_{a}(t)=x_{a}(t)\mathbf{n}+\mathbf{R}_{a}$ which represents the trajectory of the atom’s center of mass. We find it convenient to decompose the electric field operator into two contributions, i.e. $\hat{\mathbf{E}}(\mathbf{r},t)=\hat{\mathbf{E}}_{0}(\mathbf{r},t)+\hat{\mathbf{E}}_{\rm s}(\mathbf{r},t)$, where the first terms, $\hat{\mathbf{E}}_{0}(\mathbf{r},t)$, represents the electric field which would exist without the atom and we assume it to be thermalized. Consequently, this field obeys the fluctuation-dissipation theorem [3, 57, 78], and we can write
$\langle\hat{\mathbf{E}}_{0}(q;\mathbf{R}_{a};\omega)\hat{\mathbf{E}}^{\mathsf{T}}_{0}(\tilde{q},\mathbf{R}_{a};\tilde{\omega})\rangle\\ =\frac{8\pi^{2}\hbar}{1-e^{-\beta\hbar\omega}}\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\delta(\omega+\tilde{\omega})\delta(q+\tilde{q})\,,$ (30)
where $\delta(\cdot)$ is the Dirac function. The second contribution, $\hat{\mathbf{E}}_{\rm s}(\mathbf{r},t)$, describes the radiation induced by the atom. It corresponds to the special solution of Maxwell equations having the dipole as source and, as described above, it can be written as
$\hat{\mathbf{E}}_{\rm s}(\mathbf{r},t)$ $=\int\limits\mathrm{d}t^{\prime}\,\underline{G}(\mathbf{r},\mathbf{r}_{a}(t^{\prime}),t-t^{\prime})\,\hat{\mathbf{d}}(t^{\prime})~{},$ (31)
where $\underline{G}$ is the Green tensor, represented in this case in time and space.
When the NESS is achieved, we have that $x_{a}(t)\sim vt$ [57] and combining the above expression with Eq. (29), we Fourier transform into the frequency domain and can solve the system self-consistently up to all orders in the coupling $\alpha_{0}$  [78]. The stationary solution of the dipole’s dynamics takes the form
$\hat{\mathbf{d}}(\omega)=\int\frac{\mathrm{d}q}{2\pi}\underline{\alpha}(\omega,v)\cdot\hat{\mathbf{E}}_{0}(q,\mathbf{R}_{a};\omega_{q}^{+})~{},$ (32)
where $\underline{\alpha}(\omega,v)$ is the steady-state dressed atomic polarizability defined as
$\underline{\alpha}(\omega,v)$ $=\left[\underline{\mathds{1}}-\alpha_{\mathrm{B}}(\omega)\int\frac{\mathrm{d}q}{2\pi}~{}\underline{G}(q,\mathbf{R}_{a},\omega_{q}^{+})\right]^{-1}\alpha_{\mathrm{B}}(\omega)~{}.$ (33)
Here, $\alpha_{\mathrm{B}}(\omega)=\alpha_{0}\omega_{a}^{2}/\left(\omega_{a}^{2}-[\omega+\mathrm{i}0^{+}]^{2}\right)$ is the (causal) bare polarizability  [57]. We would like to note that the main difference between the bare and the dressed polarizability is the appearance of the Green tensor in Eq. (33) which indicates that the particle dynamics is modified by its interaction with the electromagnetic environment. Given that $\underline{G}^{*}(q,\mathbf{R}_{a},\omega)=\underline{G}(-q,\mathbf{R}_{a},-\omega)$, the dressed polarizability fulfills the crossing relation $\underline{\alpha}^{*}(\omega,v)=\underline{\alpha}(-\omega,v)$. Furthermore, upon utilizing the reciprocity of the involved materials, mathematically expressed by $\underline{G}^{\sf T}(q,\mathbf{R}_{a},\omega)=\underline{G}(-q,\mathbf{R}_{a},\omega)$, we can derive with the help of Eq. (33) that $\underline{\alpha}(\omega,v)=\underline{\alpha}^{\sf T}(\omega,-v)$. This result is in agreement with our general considerations. Indeed, in this case the linear polarizability does not depend on the state of the system and coincides with the expression we would have obtained from the equivalent classical system.
We are now in the position to determine the atomic power spectrum $\underline{S}(\omega,v)$ from the relation $\langle\hat{\mathbf{d}}(\omega)\hat{\mathbf{d}}^{\mathsf{T}}(\tilde{\omega})\rangle=4\pi^{2}\delta(\omega+\tilde{\omega})\underline{S}(\omega,v)$  [25, 78]. Upon considering the stationary solution for the dipole’s dynamics as given by Eq. (32) together with the properties of the polarizability and the expression for the field correlator in Eq. (31), we can write
$\underline{S}(\omega,v)$ $=\underline{\alpha}(\omega,v)\underline{\mathcal{D}}(\omega,v)\underline{\alpha}^{\dagger}(\omega,v).$ (34)
In this expression, we have introduced the tensor $\underline{\mathcal{D}}(\omega,v)$
$\underline{\mathcal{D}}(\omega,v)$ $=\frac{\hbar}{\pi}\int\frac{\mathrm{d}q}{2\pi}\frac{\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega_{q}^{+})}{1-e^{-\beta\hbar\omega^{+}_{q}}}~{}.$ (35)
With the above results, we can explicitly check the general properties of the power-spectrum tensor which we have discussed in Sec. II and at the beginning of Sec. III. To start, we notice that $\underline{S}$ is Hermitian and in general not real-valued, as expected. It does not fulfill the FDT but, instead, fulfills Eq. (13). To see this more clearly, we consider the following identity which we can derive from Eq. (33)
$\underline{\alpha}_{\Im}(\omega,v)=\int\frac{\mathrm{d}q}{2\pi}\;\underline{\alpha}(\omega,v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega_{q}^{+})\underline{\alpha}^{{\dagger}}(\omega,v)~{}.$ (36)
In fact, this equation allows us to write an explicit expression for $\underline{J}(\omega)$ as
$\underline{J}(\omega)=\frac{\hbar}{2\pi}\int\frac{{\rm d}q}{2\pi}$ $\left[\coth\left(\frac{\beta\hbar\omega^{+}_{q}}{2}\right)-\coth\left(\frac{\beta\hbar\omega}{2}\right)\right]$
×$\times\underline{\alpha}(\omega,v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega_{q}^{+})\underline{\alpha}^{{\dagger}}(\omega,v)~{}.$ (37)
Upon utilizing the properties of the polarizability and of the Green tensors mentioned above, we can show that, consistent with the discussion in Sec. II, $\underline{J}$ is Hermitian and fulfills the crossing relation. Moreover, in earlier work with our co-workers, we have demonstrated within the above framework that the local thermal equilibrium approximation displays significant drawbacks regarding the quantitative evaluation of quantum friction. Specifically, the local thermal equilibrium approximation underestimates the strength of the interaction and adversely affects the thermodynamical consistency of the physical system  [24, 25, 26].
Consider now the power-spectrum tensor. Upon noting that $(1-e^{x})^{-1}=1-(1-e^{-x})^{-1}$ and upon utilizing the expressions in Eqs. (III.1) and (36) as well as the properties of the dressed polarizability, we find
$\underline{S}(-\omega,v)=\underline{S}^{\sf T}(\omega,v)-\frac{\hbar}{\pi}\underline{\alpha}^{\sf T}_{\Im}(\omega,v)~{}.$ (38)
This coincides with Eq. (10) which we have obtained in the first part of our analysis. We may conveniently check further properties once we specify the form for the Green tensor $\underline{G}(q,\mathbf{R}_{a},\omega)$. In general, the Green tensor can be decomposed into two terms. The first term, $\underline{G}^{(\rm v)}$, contains the geometry-independent vacuum behavior and has always the same form. Due to the homogeneity and isotropy of vacuum, $\underline{G}^{(\rm v)}(q,\mathbf{R}_{a},\omega)$ does not depend on position, is symmetric and $\underline{G}_{\Im}^{(\rm v)}=\underline{G}_{I}^{(\rm v)}$ is diagonal  [100, 103]. The second term, $\underline{G}^{(\rm sc)}$, describes the scattering and absorption due to the presence of material bodies and, consequently, depends on their composition, shape and arrangement. As an example, we consider a half space, comprised of an isotropic material, located at $z<0$ and having a planar interface with vacuum at $z=0$. In this case, the mathematical expression for $\underline{G}^{(\rm sc)}$ can be found in the literature (see for example  [104, 100, 103, 57]). Due to the symmetry of the system, we may assume without loss of generality that the particle moves along the $x$-axis and choose $\mathbf{R}_{a}\equiv(y_{a}=0,z_{a})$. Indeed, since the planar structure is translationally invariant along the $y$-axis, the choice of $y_{a}=0$ is physically irrelevant. If we denote the three-dimensional wave vector with $\mathbf{K}=(k_{x},k_{y},k_{z})$, we have $q\equiv k_{x}$ in Eq. (23). In addition, we define $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$ so that $k_{z}=-\mathrm{i}\kappa=\sqrt{k^{2}-\omega^{2}/c^{2}}$ ($\mathrm{Re}[\kappa]>0$ and $\mathrm{Im}[\kappa]<0$). Based on this, we can write
$\underline{G}^{(\rm sc)}(k_{x},z_{a},\omega)=\frac{\omega^{2}}{\epsilon_{0}c^{2}}\int\frac{dk_{y}}{2\pi}\left(\underline{\Pi}_{p}r^{p}+\underline{\Pi}_{s}r^{s}\right)\frac{e^{-2\kappa z_{a}}}{2\kappa},$ (39)
where $\epsilon_{0}$ is the vacuum permittivity and $r^{\sigma}\equiv r^{\sigma}(\omega,k)$ denote the polarization-dependent ($\sigma=s,p$) reflection coefficients of the surface [105]. The expressions for $r^{\sigma}$ depend on the specific properties of the material half-space but, due to its isotropy, they can only depend on $k$. Furthermore, in the above expressions, we have defined the polarization matrices $\underline{\Pi}_{p}$ and $\underline{\Pi}_{s}$, which, owing to the symmetries of the expression in Eq. (39), can be written as
$\underline{\Pi}_{s}=\begin{pmatrix}\frac{k_{y}^{2}}{k^{2}}&0&0\\ 0&\frac{k_{x}^{2}}{k^{2}}&0\\ 0&0&0\end{pmatrix},\quad\underline{\Pi}_{p}=\frac{c^{2}\kappa^{2}}{\omega^{2}}\begin{pmatrix}\frac{k_{x}^{2}}{k^{2}}&0&-\mathrm{i}\frac{k}{\kappa}\frac{k_{x}}{k}\\ 0&\frac{k_{y}^{2}}{k^{2}}&0\\ \mathrm{i}\frac{k}{\kappa}\frac{k_{x}}{k}&0&\frac{k^{2}}{\kappa^{2}}\end{pmatrix}.$ (46)
With the above, we can already check that the expression in Eq. (39) has a form that is consistent with Eq. (26) where $c_{n}=0$ and $\mathbf{s}_{\perp}=s_{y}\mathbf{y}$. As discussed earlier, we may view this as being the consequence of the mirror-symmetry of the total system with respect to the $y=0$ plane. Finally, inserting in the previous equations the expression for the Green tensor given above and using Eq. (34), one can show that $\underline{S}(\omega,v)$ is given by the sum of a diagonal and an anti-diagonal matrix. This is in agreement with our general considerations about the structure of the tensor based on the symmetry of the system. Moreover, we can verify that $\underline{S}(\omega,-v)\neq\underline{S}^{\sf T}(\omega,v)$, as we should expect in terms of our general result in Eq. (20). Notice that this result is obtained despite the fact that $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$ and that from Eq. (35) $\underline{\mathcal{D}}(\omega,-v)=\underline{\mathcal{D}}^{\sf T}(\omega,v)$. One can verify that
$\underline{S}(\omega,-v)-\underline{S}^{\sf T}(\omega,v)=\begin{pmatrix}0&0&f(\omega,v)\\ 0&0&0\\ f(\omega,v)&0&0\end{pmatrix}$ (47)
where, in agreement with the relation in Eq. (20), $f(\omega,v)$ is a real function odd in $v$. The previous expression also implicitly confirms that the diagonal elements of the power-spectrum tensor are even in $v$, while the off-diagonal terms arising from $\bm{\Omega}_{\perp}=\Omega_{y}\mathbf{y}$ are odd in the same variable.
IV Conclusions
In summary, we have discussed a number of properties of the power-spectral density for a quantum system in an arbitrary steady state. This includes a large class of nonequilibrium scenarios as well as the equilibrium situation as a limiting case. We have adopted a minimalistic viewpoint and, besides stationarity, our analyses do not rely on any specific information about the microscopic dynamics of the system. Therefore, our results have a broad validity, including systems with non-Markorvian as well as nonlinear dynamics. This is particularly relevant for quantum systems in which the Markov approximation may lead to inconsistent results [4, 57].
More precisely, we have focused on the steady state of a multidimensional (multivariate) stochastic process, where the dynamics is described in terms of a three-dimensional vector operator. Since the steady state may correspond to nonequilibrium dynamics for which the principle of detailed balance can be violated, equilibrium relations such as the fluctuation-dissipation theorem are not necessarily fulfilled. Nonetheless, we have been able to derive a collection of properties that characterize the power spectrum and its connection to other physical quantities such as the system’s susceptibility. In particular, we have determined the properties of quantities which can characterize the system’s steady-state nonequilibrium behavior and have defined the correction to the equilibrium fluctuation-dissipation theorem. This provides us with a solid basis to investigate and assess the validity of approaches such as the local thermal equilibrium approximation.
Once additional information about the system such as symmetries and details about its structure is provided, we may obtain further relations and properties among the different quantities that characterize the system. As an illustration, we have analyzed the nonconservative viscous dynamics of a particle that moves in the quantum electromagnetic vacuum parallel to a translationally invariant arrangement of material bodies. When the drag felt by the particle during its motion is balanced by an external driving force, the system reaches a nonequilibrium steady state that is characterized by a speed $v$. Firstly, without specifying the internal microscopic dynamics of the moving object, we have derived a collection of relations that characterize the power-spectrum tensor, its dependence on the velocity $v$ and its role in determining the frictional interaction at zero temperature (quantum friction). Secondly, for this system, we have tested the general results described in main part of the manuscript, in the context of a specific model which provides the microscopic internal dynamics of a moving atom. In particular, we have modeled the atom in terms of a three-dimensional harmonic oscillator and have derived and examined the corresponding power-spectrum tensor as well as the related matrices.
Our results can be useful in a large variety of systems and specifically for numerous fluctuation-induced phenomena, such as the (quantum) Brownian motion, the Casimir effect and the van der Waals force, including their nonequilbrium variants as quantum friction. Since often the information about the statistical properties of the system is or can be encoded in the power spectral density, knowing its general properties can be very helpful for understanding its behavior and to test thermodynamic consistency of its descriptions [26].
V Acknowledgements
We sincerely thank Bettina Beverungen, Daniel Reiche and Marty Oelschläger for numerous motivating questions and valuable discussions around many of the results presented here.
Appendix A Positive definite matrices
In this appendix we elaborate on the reasons for why some of the matrices encountered in the main text are positive semidefinite or positive definite.
Let us consider first the power-spectrum tensor. For a generic complex vector $\mathbf{u}$ and any real-valued function $f(t)$ we can define the operator
[…]
Since $\langle\hat{b}^{{\dagger}}_{\mathbf{u}}(t)\hat{b}_{\mathbf{u}}(t)\rangle$ is always nonnegative it follows that
[…]
Then, since $|f(\omega)|^{2}\geq 0$, one must have that
$\mathbf{u}^{*}\cdot\underline{S}(\omega)\cdot\mathbf{u}\geq 0\quad\forall\mathbf{u}~{}\forall\omega~{},$ (50)
which proves that $\underline{S}(\omega)$ is positive semidefinite.
For discussing the properties of $\underline{\alpha}_{\Im}(\omega)$ we consider the perturbed Hamiltonian $\hat{H}^{\prime}_{t}=\hat{H}_{0}-\hat{\mathbf{d}}\cdot\mathbf{e}(t)$, where $\mathbf{e}(t)$ denotes a small ($\left|\mathbf{e}(t)\right|\ll 1$) real three-dimensional vector which is different from zero for a finite interval of time $\mathbf{e}(t\to\pm\infty)\to 0$. In linear order of perturbation theory, the expectation value of the associated time evolution of the dipole operator as prescribed by $\hat{H}^{\prime}_{t}$, i.e. $\hat{\mathbf{d}}^{\prime}(t)$, can be written as
$\langle\hat{\mathbf{d}}^{\prime}(t)\rangle\equiv\mathrm{tr}[\hat{\mathbf{d}}^{\prime}(t)\hat{\rho}]$ $\stackrel{{\scriptstyle\left|\mathbf{e}(t)\right|\ll 1}}{{\approx}}\int_{-\infty}^{\infty}\mathrm{d}t_{1}\;\underline{\alpha}(t-t_{1})\cdot\mathbf{e}(t_{1})~{},$ (51)
where $\underline{\alpha}(\tau)$ is the linear response tensor with elements defined in Eq. (6). Mathematically, we may regard the tensor as the functional derivative with respect to the external perturbation of the exact time evolution, i.e. $\delta\langle\hat{d}^{\prime}_{i}(t)\rangle/\delta e_{j}(t^{\prime})=\alpha_{ij}(t-t^{\prime})$.
In general, the time evolution according to $\hat{H}^{\prime}_{t}$ of an operator $\hat{A}_{t}$ which explicitly depends on time is given by $\hat{A}_{t}(t)=U^{{\dagger}}_{t}\hat{A}_{t}U_{t}$,. The operator $U_{t}$ is the solution of $\partial_{t}U_{t}=(\mathrm{i}/\hbar)\hat{H}_{t}\hat{U}_{t}$ with the initial condition $U_{t=0}=\underline{\mathds{1}}$. Accordingly, the total time derivative of $\hat{A}_{t}(t)$ is given by
$\frac{d}{dt}\hat{A}_{t}(t)=\frac{\mathrm{i}}{\hbar}U^{{\dagger}}_{t}[\hat{H}_{t},\hat{A}_{t}]U_{t}+U^{{\dagger}}_{t}(\partial_{t}\hat{A}_{t})U_{t}~{}.$ (52)
As a consequence, during the perturbation, the total change of energy is given by
[…]
$=-\int_{-\infty}^{\infty}\mathrm{d}t\;\langle\hat{\mathbf{d}}^{\prime}(t)\rangle\cdot\frac{d\mathbf{e}(t)}{dt}$ (53)
Inserting Eq. (51) we have that
[…]
where $\underline{\alpha}_{\Im}(\omega)$ is the the matrix defined in Eq. (7) of the main text. If we require that the energy dissipated during the perturbation is always positive irrespective of the form of the perturbation, then for $\omega>0$ the matrix $\underline{\alpha}_{\Im}(\omega)$ must be positive definite.
Finally, the previous considerations allow us to say that the matrix $\underline{\nu}(\omega)$ defined in Eq. (9) is also positive semidefinite [32]. Indeed, using its definition and properties discussed in the main text, we have that
$\underline{\nu}(\omega)=\underline{\nu}^{*}(-\omega)=\underline{S}^{*}(-\omega)+\frac{\hbar}{2\pi}\underline{\alpha}_{\Im}(\omega)~{},$ (55)
or also combining this again with Eq. (9) that
$\underline{\nu}(\omega)=\frac{\underline{S}(\omega)+\underline{S}^{*}(-\omega)}{2}=\frac{\underline{S}(\omega)+\underline{S}^{\sf T}(-\omega)}{2}~{}.$ (56)
Both expression show that $\underline{\nu}(\omega)$ for $\omega\geq 0$ is positive semidefinite because it is given by the sum of positive semidefinite matrices.
Appendix B Symmetry properties
In this appendix we provide additional information about the consequences of the mirror-symmetry on the frictional quantum system discussed in Sec. III.
We start by considering the tensor $\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)$. Since it is Hermitian, we can write
$\underline{G}_{\Im}$ $(q,\mathbf{R}_{a},\omega)=\underline{D}(q,\mathbf{R}_{a},\omega)$
$+\mathbf{c}(q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{\Lambda}}+\mathbf{s}(q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{L}}$
$=\underline{D}(q,\mathbf{R}_{a},\omega)$
$+c_{n}(q,\mathbf{R}_{a},\omega)\underline{\Lambda}_{n}+\mathbf{c}_{\perp}(q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$+s_{n}(q,\mathbf{R}_{a},\omega)\underline{L}_{n}+\mathbf{s}_{\perp}(q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{L}}_{\perp}~{},$ (57)
where $\underline{D}(q,\mathbf{R}_{a},\omega)$ is a diagonal matrix, $\mathbf{c}(q,\mathbf{R}_{a},\omega)$ and $\mathbf{s}(q,\mathbf{R}_{a},\omega)$ are two real-valued three-dimensional vectors. The matrices $\underline{\Lambda}_{i}$ and $\underline{L}_{i}$ are defined in the main text (see the text below Eqs. (4) and (19)). The subscript $n$ and $\perp$, respectively, indicate the components parallel and orthogonal to the direction of translational invariance that is given by $\mathbf{n}$. Since this decomposition is unique, we can utilize Eq. (24a) and the properties of the involved matrices to conclude that $\underline{D}$ and $\mathbf{c}$ are even functions of $q$, while $\mathbf{s}$ is an odd function of $q$.
If we consider the symmetry described by $\underline{M}_{n}$, Eq. (25) reveals that for the setup under consideration the above expression must coincide with
$\underline{M}_{n}\underline{G}_{\Im}$ $(-q,\mathbf{R}_{a},\omega)\underline{M}^{-1}_{n}=\underline{D}(-q,\mathbf{R}_{a},\omega)$
$+c_{n}(-q,\mathbf{R}_{a},\omega)\underline{\Lambda}_{n}-\mathbf{c}_{\perp}(-q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$+s_{n}(-q,\mathbf{R}_{a},\omega)\underline{L}_{n}-\mathbf{s}_{\perp}(-q,\mathbf{R}_{a},\omega)\cdot\mathbf{\underline{L}}_{\perp}~{}.$ (58)
Here, we have used that diagonal matrices and $\underline{\Lambda}_{n}$ are invariant with respect to the transformation while for the other matrices we have $\underline{M}_{n}\bm{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\bm{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. When comparing the coefficients of Eqs. (B) and (B), we obtain $\mathbf{c}_{\perp}(-q,\mathbf{R}_{a},\omega)=-\mathbf{c}_{\perp}(q,\mathbf{R}_{a},\omega)$ (odd function in $q$) and $s_{n}(q,\mathbf{R}_{a},\omega)=s_{n}(-q,\mathbf{R}_{a},\omega)$ (even function in $q$). This is, however, in contradiction with the the constraints given by Eq. (24a) and, consequently, these expressions must vanish. This leads to the decomposition in Eq. (26). We can follow a similar procedure if the system features an additional symmetry $\underline{M}_{t}$, where $\mathbf{t}\cdot\mathbf{n}=0$. If the symmetry plane contains both, $\mathbf{q}=q\mathbf{n}$ and $\mathbf{R}_{a}$, thus making them invariant with respect to $\underline{M}_{t}$, we must have that $\underline{G}_{\Im}$ in Eq. (26) is invariant with respect to this transformation. However, if we call $\mathbf{t^{\prime}}=\mathbf{n}\times\mathbf{t}$, we obtain that
$\underline{G}_{\Im}=\underline{D}+c_{n}\underline{\Lambda}_{n}+s_{t^{\prime}}\underline{L}_{t^{\prime}}+s_{t}\underline{L}_{t}$
$=\underline{M}_{t}\underline{G}_{\Im}\underline{M}_{t}^{-1}=\underline{D}-c_{n}\underline{\Lambda}_{n}-s_{t^{\prime}}\underline{L}_{t^{\prime}}+s_{t}\underline{L}_{t},$ (59)
which, in turn, implies that $c_{n}=s_{t^{\prime}}=0$.
The power-spectrum tensor $\underline{S}(\omega,v)$ is subject to fewer restrictions. In general, we have that
$\underline{S}(\omega,v)$ $=\underline{\Delta}(\omega,v)+\bm{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}+\bm{\Omega}(\omega,v)\cdot\mathbf{\underline{L}}$
$=\underline{\Delta}(\omega,v)$
$\hskip 28.45274pt+\chi_{n}(\omega,v)\underline{\Lambda}_{n}+\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$\hskip 28.45274pt+\Omega_{n}(\omega,v)\underline{L}_{n}+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~{}.$ (60)
Due to the symmetry of the system and its dissipative nature, we can anticipate that $\Omega_{n}=0$. Arguments analogous to those made for $\underline{G}_{\Im}$ reveal that $\underline{\Delta}$ and $\chi_{n}$ are even functions of the velocity $v$, while $\bm{\chi}_{\perp}$ and $\bm{\Omega}_{\perp}$ must be a odd function of $v$. Therefore, we can write
$\underline{S}^{\sf T}(\omega,-v)$ $=\underline{\Delta}^{\sf T}(\omega,-v)+\bm{\chi}(\omega,-v)\cdot\mathbf{\underline{\Lambda}}^{\sf T}$
$\hskip 59.75095pt+\bm{\Omega}_{\perp}(\omega,-v)\cdot\mathbf{\underline{L}}_{\perp}^{\sf T}$
$=\underline{\Delta}(\omega,v)+\chi_{n}(\omega,v)\underline{\Lambda}_{n}-\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$\hskip 59.75095pt+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}$
$=\underline{S}(\omega,v)-2\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~{},$ (61)
which is equivalent to Eq. (20). Again, when an additional mirror-symmetry $\underline{M}_{t}$ occurs and $\mathbf{v}=v\mathbf{n}$ and $\mathbf{R}_{a}$ are contained in (or parallel to) the symmetry plane, then $\underline{S}(\omega,v)$ must be invariant with respect to the transformation. With arguments analogous to the above-discussed case of $\underline{G}_{\Im}$, we can then conclude that $\bm{\chi}=\chi_{t}\mathbf{t}$ and $\bm{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$.